Find the Cartesian equation of r = acosθ/2
Answer by stanbon(75887)

Find the Cartesian equation of r = acosθ/2
------
r = sqrt(x^2+y^2)
theta = arctan(y/x)
---
Since tan(theta) = y/x, y = sin(theta) and x = cos(theta)
------------------------------------
Therefore cos(theta/2) = sqrt[(1+cos(theta))/2] = sqrt[1+(x/2)]
----
Equation:
x^2 + y^2 = a*sqrt(1+(x+2)]
------------
Cheers,
Stan H.
-------------